deathscarewolf:

again comprehensive science During an investigation a scientist heated 2.76 g of silver carbonate till it decomposed to leave a silver residue. The total mass of the silver residue formed was 2.16 g. Does the law of conservation of mass hold true in this case? Use complete sentences to justify your answer based on numerical calculations.

SweetPumpkin1775:

The equation that shows the decomposition of silver carbonate is as follows: 2Ag2CO3 .........> 4Ag + 2CO2 + O2 From the periodic table: mass of silver = 107.8682 grams mass of carbon = 12 grams mass of oxygen = 16 grams molar mass of Ag2CO3 = 2(107.8682) + 12 + 3(16) = 275.7364 grams From the balanced equation above: 2(275.7362) = 551.4728 grams of Ag2CO3 produces 4(107.8682) = 431.4728 grams of Ag Therefore, to know the mass of Ag produced from 2.76 grams of Ag2CO3, we will simply use cross multiplication as follows: mass of Ag produced = (2.76*431.4728) / (551.4728) = 2.16 grams Based on the above calculations, the law of conservation of mass is applied. Muxakara and 85 more users found this answer helpful THANKS 47 4.0 (38 votes) Need more info? Try these questions burtonr080504 avatar burtonr080504 02/26/2018 Physics High School answered • expert verified During an investigation a scientist heated 2.76 g of silver carbonate till it decomposed to leave a silver residue. The total mass of the silver residue formed was 2.16 g. Does the law of conservation of mass hold true in this case? Use complete sentences to justify your answer based on numerical calculations. See answers (1) homeschool0123 avatar homeschool0123 11/06/2018 Biology High School answered LAST ONE!!! During an investigation a scientist heated 2.76 g of silver carbonate till it decomposed to leave a silver residue. The total mass of the silver residue formed was 2.16 g. Does the law of conservation of mass hold true in this case? Use complete sentences to justify your answer based on numerical calculations. See answers (1) chinared221981 10/11/2017 Physics Middle School answered During an investigation, a scientist heated 123.6 g of copper carbonate till it decomposed to form a black residue. The total mass of the black residue formed was 79.6 g. Does the law of conservation of mass hold true in this case? Use complete sentences to justify your answer based on numerical calculations. See answers (1) Still have questions? FIND MORE ANSWERS ASK YOUR QUESTION New questions in Chemistry Storms are blank weather In what way are the two ways of treating sewage really just ways of recycling? Do any body wants to role play Notes: Thermal Energy Instructions Read the webpage on Thermal Energy Watch the video at the end of the notes Upload your answers to the Review Qu… plants get energy from the​ Which of the following is a physical or chemical characteristic of a mineral? im trying to understand how to solve chemical problem​ like mg + o2 ---->mgo Why is human waste hazardous when it enters the water system? for London and Ancient Rome Pls Help don't give me no 5-year old answers of"haha free points hehe"cuz it an no yi ng so pls help the diagram below shows a food web. If a certai… would you want to build the wire for a cell phone charger out of nitrogen (and) why or why not Question 13 1 pts A solution is formed by dissolving 4.81g of sodium sulfate, Na2SO4, into 0.210L of water. What is the volume of the solution for… In covalent compounds the bonding (valence) electrons are Hint: The answer is a verb that ends with red. Hint: Think of co-champs. I’m sad and low yeah I’m sad and low yeahhh the big bored........XD Find the cube of each semimajor axis length (A) by raising the value to the third power. Write your results in the table provided. Round all values t… The nucleus of an atom is made up of a dense ball of_________and_____________surrounded by________. The reaction A(g) ⇌ 2 B(g) is exothermic. Identify the direction of the shift in equilibrium after each of the following changes: a. Increase [A] __… Identify the phase change described by the following equations. a. C18H38(s) → C18H38(l) _______________________ b. CO2(g) → CO2(s) ______________… Which of the following statements about the sample of Moscovium is correct? How many moles of aluminum oxide will be formed from 17 moles of aluminum reaction? __Al + __ O2 — __Al2O3 Quoted from a cited called Brainly.... Qc is better tho.

deathscarewolf:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @SweetPumpkin1775 The equation that shows the decomposition of silver carbonate is as follows: 2Ag2CO3 .........> 4Ag + 2CO2 + O2 From the periodic table: mass of silver = 107.8682 grams mass of carbon = 12 grams mass of oxygen = 16 grams molar mass of Ag2CO3 = 2(107.8682) + 12 + 3(16) = 275.7364 grams From the balanced equation above: 2(275.7362) = 551.4728 grams of Ag2CO3 produces 4(107.8682) = 431.4728 grams of Ag Therefore, to know the mass of Ag produced from 2.76 grams of Ag2CO3, we will simply use cross multiplication as follows: mass of Ag produced = (2.76*431.4728) / (551.4728) = 2.16 grams Based on the above calculations, the law of conservation of mass is applied. Muxakara and 85 more users found this answer helpful THANKS 47 4.0 (38 votes) Need more info? Try these questions burtonr080504 avatar burtonr080504 02/26/2018 Physics High School answered • expert verified During an investigation a scientist heated 2.76 g of silver carbonate till it decomposed to leave a silver residue. The total mass of the silver residue formed was 2.16 g. Does the law of conservation of mass hold true in this case? Use complete sentences to justify your answer based on numerical calculations. See answers (1) homeschool0123 avatar homeschool0123 11/06/2018 Biology High School answered LAST ONE!!! During an investigation a scientist heated 2.76 g of silver carbonate till it decomposed to leave a silver residue. The total mass of the silver residue formed was 2.16 g. Does the law of conservation of mass hold true in this case? Use complete sentences to justify your answer based on numerical calculations. See answers (1) chinared221981 10/11/2017 Physics Middle School answered During an investigation, a scientist heated 123.6 g of copper carbonate till it decomposed to form a black residue. The total mass of the black residue formed was 79.6 g. Does the law of conservation of mass hold true in this case? Use complete sentences to justify your answer based on numerical calculations. See answers (1) Still have questions? FIND MORE ANSWERS ASK YOUR QUESTION New questions in Chemistry Storms are blank weather In what way are the two ways of treating sewage really just ways of recycling? Do any body wants to role play Notes: Thermal Energy Instructions Read the webpage on Thermal Energy Watch the video at the end of the notes Upload your answers to the Review Qu… plants get energy from the​ Which of the following is a physical or chemical characteristic of a mineral? im trying to understand how to solve chemical problem​ like mg + o2 ---->mgo Why is human waste hazardous when it enters the water system? for London and Ancient Rome Pls Help don't give me no 5-year old answers of"haha free points hehe"cuz it an no yi ng so pls help the diagram below shows a food web. If a certai… would you want to build the wire for a cell phone charger out of nitrogen (and) why or why not Question 13 1 pts A solution is formed by dissolving 4.81g of sodium sulfate, Na2SO4, into 0.210L of water. What is the volume of the solution for… In covalent compounds the bonding (valence) electrons are Hint: The answer is a verb that ends with red. Hint: Think of co-champs. I’m sad and low yeah I’m sad and low yeahhh the big bored........XD Find the cube of each semimajor axis length (A) by raising the value to the third power. Write your results in the table provided. Round all values t… The nucleus of an atom is made up of a dense ball of_________and_____________surrounded by________. The reaction A(g) ⇌ 2 B(g) is exothermic. Identify the direction of the shift in equilibrium after each of the following changes: a. Increase [A] __… Identify the phase change described by the following equations. a. C18H38(s) → C18H38(l) _______________________ b. CO2(g) → CO2(s) ______________… Which of the following statements about the sample of Moscovium is correct? How many moles of aluminum oxide will be formed from 17 moles of aluminum reaction? __Al + __ O2 — __Al2O3 Quoted from a cited called Brainly.... Qc is better tho. \(\color{#0cbb34}{\text{End of Quote}}\) some cant get onto brainly anymore

SweetPumpkin1775:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @deathscarewolf \(\color{#0cbb34}{\text{Originally Posted by}}\) @SweetPumpkin1775 The equation that shows the decomposition of silver carbonate is as follows: 2Ag2CO3 .........> 4Ag + 2CO2 + O2 From the periodic table: mass of silver = 107.8682 grams mass of carbon = 12 grams mass of oxygen = 16 grams molar mass of Ag2CO3 = 2(107.8682) + 12 + 3(16) = 275.7364 grams From the balanced equation above: 2(275.7362) = 551.4728 grams of Ag2CO3 produces 4(107.8682) = 431.4728 grams of Ag Therefore, to know the mass of Ag produced from 2.76 grams of Ag2CO3, we will simply use cross multiplication as follows: mass of Ag produced = (2.76*431.4728) / (551.4728) = 2.16 grams Based on the above calculations, the law of conservation of mass is applied. Muxakara and 85 more users found this answer helpful THANKS 47 4.0 (38 votes) Need more info? Try these questions burtonr080504 avatar burtonr080504 02/26/2018 Physics High School answered • expert verified During an investigation a scientist heated 2.76 g of silver carbonate till it decomposed to leave a silver residue. The total mass of the silver residue formed was 2.16 g. Does the law of conservation of mass hold true in this case? Use complete sentences to justify your answer based on numerical calculations. See answers (1) homeschool0123 avatar homeschool0123 11/06/2018 Biology High School answered LAST ONE!!! During an investigation a scientist heated 2.76 g of silver carbonate till it decomposed to leave a silver residue. The total mass of the silver residue formed was 2.16 g. Does the law of conservation of mass hold true in this case? Use complete sentences to justify your answer based on numerical calculations. See answers (1) chinared221981 10/11/2017 Physics Middle School answered During an investigation, a scientist heated 123.6 g of copper carbonate till it decomposed to form a black residue. The total mass of the black residue formed was 79.6 g. Does the law of conservation of mass hold true in this case? Use complete sentences to justify your answer based on numerical calculations. See answers (1) Still have questions? FIND MORE ANSWERS ASK YOUR QUESTION New questions in Chemistry Storms are blank weather In what way are the two ways of treating sewage really just ways of recycling? Do any body wants to role play Notes: Thermal Energy Instructions Read the webpage on Thermal Energy Watch the video at the end of the notes Upload your answers to the Review Qu… plants get energy from the​ Which of the following is a physical or chemical characteristic of a mineral? im trying to understand how to solve chemical problem​ like mg + o2 ---->mgo Why is human waste hazardous when it enters the water system? for London and Ancient Rome Pls Help don't give me no 5-year old answers of"haha free points hehe"cuz it an no yi ng so pls help the diagram below shows a food web. If a certai… would you want to build the wire for a cell phone charger out of nitrogen (and) why or why not Question 13 1 pts A solution is formed by dissolving 4.81g of sodium sulfate, Na2SO4, into 0.210L of water. What is the volume of the solution for… In covalent compounds the bonding (valence) electrons are Hint: The answer is a verb that ends with red. Hint: Think of co-champs. I’m sad and low yeah I’m sad and low yeahhh the big bored........XD Find the cube of each semimajor axis length (A) by raising the value to the third power. Write your results in the table provided. Round all values t… The nucleus of an atom is made up of a dense ball of_________and_____________surrounded by________. The reaction A(g) ⇌ 2 B(g) is exothermic. Identify the direction of the shift in equilibrium after each of the following changes: a. Increase [A] __… Identify the phase change described by the following equations. a. C18H38(s) → C18H38(l) _______________________ b. CO2(g) → CO2(s) ______________… Which of the following statements about the sample of Moscovium is correct? How many moles of aluminum oxide will be formed from 17 moles of aluminum reaction? __Al + __ O2 — __Al2O3 Quoted from a cited called Brainly.... Qc is better tho. \(\color{#0cbb34}{\text{End of Quote}}\) some cant get onto brainly anymore \(\color{#0cbb34}{\text{End of Quote}}\) hmm?

deathscarewolf:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @SweetPumpkin1775 \(\color{#0cbb34}{\text{Originally Posted by}}\) @deathscarewolf \(\color{#0cbb34}{\text{Originally Posted by}}\) @SweetPumpkin1775 The equation that shows the decomposition of silver carbonate is as follows: 2Ag2CO3 .........> 4Ag + 2CO2 + O2 From the periodic table: mass of silver = 107.8682 grams mass of carbon = 12 grams mass of oxygen = 16 grams molar mass of Ag2CO3 = 2(107.8682) + 12 + 3(16) = 275.7364 grams From the balanced equation above: 2(275.7362) = 551.4728 grams of Ag2CO3 produces 4(107.8682) = 431.4728 grams of Ag Therefore, to know the mass of Ag produced from 2.76 grams of Ag2CO3, we will simply use cross multiplication as follows: mass of Ag produced = (2.76*431.4728) / (551.4728) = 2.16 grams Based on the above calculations, the law of conservation of mass is applied. Muxakara and 85 more users found this answer helpful THANKS 47 4.0 (38 votes) Need more info? Try these questions burtonr080504 avatar burtonr080504 02/26/2018 Physics High School answered • expert verified During an investigation a scientist heated 2.76 g of silver carbonate till it decomposed to leave a silver residue. The total mass of the silver residue formed was 2.16 g. Does the law of conservation of mass hold true in this case? Use complete sentences to justify your answer based on numerical calculations. See answers (1) homeschool0123 avatar homeschool0123 11/06/2018 Biology High School answered LAST ONE!!! During an investigation a scientist heated 2.76 g of silver carbonate till it decomposed to leave a silver residue. The total mass of the silver residue formed was 2.16 g. Does the law of conservation of mass hold true in this case? Use complete sentences to justify your answer based on numerical calculations. See answers (1) chinared221981 10/11/2017 Physics Middle School answered During an investigation, a scientist heated 123.6 g of copper carbonate till it decomposed to form a black residue. The total mass of the black residue formed was 79.6 g. Does the law of conservation of mass hold true in this case? Use complete sentences to justify your answer based on numerical calculations. See answers (1) Still have questions? FIND MORE ANSWERS ASK YOUR QUESTION New questions in Chemistry Storms are blank weather In what way are the two ways of treating sewage really just ways of recycling? Do any body wants to role play Notes: Thermal Energy Instructions Read the webpage on Thermal Energy Watch the video at the end of the notes Upload your answers to the Review Qu… plants get energy from the​ Which of the following is a physical or chemical characteristic of a mineral? im trying to understand how to solve chemical problem​ like mg + o2 ---->mgo Why is human waste hazardous when it enters the water system? for London and Ancient Rome Pls Help don't give me no 5-year old answers of"haha free points hehe"cuz it an no yi ng so pls help the diagram below shows a food web. If a certai… would you want to build the wire for a cell phone charger out of nitrogen (and) why or why not Question 13 1 pts A solution is formed by dissolving 4.81g of sodium sulfate, Na2SO4, into 0.210L of water. What is the volume of the solution for… In covalent compounds the bonding (valence) electrons are Hint: The answer is a verb that ends with red. Hint: Think of co-champs. I’m sad and low yeah I’m sad and low yeahhh the big bored........XD Find the cube of each semimajor axis length (A) by raising the value to the third power. Write your results in the table provided. Round all values t… The nucleus of an atom is made up of a dense ball of_________and_____________surrounded by________. The reaction A(g) ⇌ 2 B(g) is exothermic. Identify the direction of the shift in equilibrium after each of the following changes: a. Increase [A] __… Identify the phase change described by the following equations. a. C18H38(s) → C18H38(l) _______________________ b. CO2(g) → CO2(s) ______________… Which of the following statements about the sample of Moscovium is correct? How many moles of aluminum oxide will be formed from 17 moles of aluminum reaction? __Al + __ O2 — __Al2O3 Quoted from a cited called Brainly.... Qc is better tho. \(\color{#0cbb34}{\text{End of Quote}}\) some cant get onto brainly anymore \(\color{#0cbb34}{\text{End of Quote}}\) hmm? \(\color{#0cbb34}{\text{End of Quote}}\) i cant get on to brainly from stupid school block

SweetPumpkin1775:

Oh ok. well.... heres the answer The equation that shows the decomposition of silver carbonate is as follows: 2Ag2CO3 .........> 4Ag + 2CO2 + O2 From the periodic table: mass of silver = 107.8682 grams mass of carbon = 12 grams mass of oxygen = 16 grams molar mass of Ag2CO3 = 2(107.8682) + 12 + 3(16) = 275.7364 grams From the balanced equation above: 2(275.7362) = 551.4728 grams of Ag2CO3 produces 4(107.8682) = 431.4728 grams of Ag Therefore, to know the mass of Ag produced from 2.76 grams of Ag2CO3, we will simply use cross multiplication as follows: mass of Ag produced = (2.76*431.4728) / (551.4728) = 2.16 grams Based on the above calculations, the law of conservation of mass is applied.

deathscarewolf:

thx